# Statistical inference on partially shape-constrained function-on-scalar linear regression models

### v1

@@ Abstract
We consider functional linear regression models where functional outcomes are associated with scalar predictors by coefficient functions with shape constraints, such as monotonicity and convexity, that apply to sub-domains of interest. To validate the partial shape constraints, we propose testing a composite hypothesis of linear functional constraints on regression coefficients. Our approach employs kernel- and spline-based methods within a unified inferential framework, evaluating the statistical significance of the hypothesis by measuring an $L^{2}$-distance between constrained and unconstrained model fits. In the theoretical study of large-sample analysis under mild conditions, we show that both methods achieve the standard rate of convergence observed in the nonparametric estimation literature. Through numerical experiments of finite-sample analysis, we demonstrate that the type I error rate keeps the significance level as specified across various scenarios and that the power increases with sample size, confirming the consistency of the test procedure under both estimation methods. Our theoretical and numerical results provide researchers the flexibility to choose a method based on computational preference. The practicality of partial shape-constrained inference is illustrated by two data applications: one involving clinical trials of NeuroBloc in type A-resistant cervical dystonia and the other with the National Institute of Mental Health Schizophrenia Study.
Keywords: Nonparametric estimation, partial shape constraints, shape-constrained kernel least squares, shape-constrained regression spline, testing
@@ 1 Introduction
As function-valued data acquisition becomes increasingly common, there has been a significant amount of work devoted to functional regression models to address coefficient function estimation and inferential capabilities [21, 32]. Among them, inference for the function-on-scalar regression (FoSR) model has been making implications in many fields [24] by identifying a dynamic association between response and scalar covariates, varying over the domain. Besides finding the statistical evidence of a non-null covariate effect on response trajectories over the domain, validating the shape of functional regression coefficients, such as monotonicity, convexity, or concavity, over a specific subset of the domain is crucial for domain experts to allow tangible interpretation in practice. Once the functional shape is validated, one may incorporate the corresponding conditions to the model estimates to avoid potential biases. Or, in practical estimation problems, shape restrictions on functional regression coefficients over specific sub-intervals of the domain can be known as prior knowledge.
Our motivating example is the functional data analysis approach to demonstrating the treatment efficacy of drug use during the period of clinical trials in the National Institute of Mental Health Schizophrenia Collaborative Study. The previous studies, including [1] and [12], showed a significant continual drop in disease severity for the treatment group, i.e., improving drug efficacy, over weeks through the shape-constraint test on the drug effect coefficient function. While the study design and models are provided in Section 4.2, the left panel of Figure 1 displays the collected longitudinal disease severity measures surveyed from the initial treatment (week 0) to week 6 from randomly selected 30 patients of placebo and treatment cohorts. However, refined conclusions over specific periods can be of practical interest to practitioners, such as when the maximum drug efficacy is achieved or whether we can conclude significant improvement in drug effectiveness even during the later weeks of trials. As we shall see in Section 4.2, our proposed partial inference method concludes the significant monotone decrease in disease severity owing to the drug treatment over weeks 0–3, followed by the consistent duration of such efficacy for the remainder weeks, illustrated with the blue solid line in the right panel of Figure 1. This is somewhat distinct from constrained estimates under the condition of monotone decrease over the entire domain, the conclusion of [12], and unconstrained smooth estimates, which are displayed with the black dotted and dashed lines, respectively.
@@
Figure 1: The left panel illustrates the individual trajectories of disease severity for a subset of patients in the treatment and placebo groups. The average disease severity for each group (solid line) shows an overall decrease, with the treatment group experiencing a rapid drop in the first few weeks, followed by a continual reduction at a seemingly similar rate for both groups. As shown in the right panel, [12] previously validated this observation using globally monotone shape-constrained estimates (black dashed line), which provide more explicit and decision-making interpretations than unconstrained estimates (black dotted line). However, the global shape-constrained inference may be misleading by ignoring the rebound after Week 5, as indicated by the unconstrained estimates. Our proposed method (blue solid line) with partial shape constraints refines the conclusion, suggesting that the duration of drug action may effectively extend up to Week 3, with sustained efficacy thereafter.
@@
The recent literature on functional data analysis and its applications also pays attention to shape-restricted estimation and inference in functional regression models. For the inference on shape-constrained regression coefficients, [6] proposed the goodness-of-fit test based on empirical processes projected to the space with shape constraints, while [22] addresses a similar problem under the null hypothesis with linear operator constraints. For estimation, [4] extended the shape-constrained kernel smoothing to the functional and longitudinal data framework, and [12] employed the Bernstein polynomials for shape-constrained estimation in functional regression with one of the data application results from [12] being illustrated in Figure 1. To name of few other applications, in economics, motivated by the pioneering work from Slutsky [31] recognizing the necessity of shape-restricted estimation and inference, [5] elaborated theoretical and application works in econometric research. In public health research, the analysis of growth charts under the monotone increasing shape restrictions has provided a crucial clinical tool for growth screening during infancy, childhood, and adolescence [15] or such growth chart helps health providers assess monotone increasing growth patterns against age-specific percentile curves [9]. The reliability engineering society also paid attention to this topic for bathtub-shaped function estimation in assessing the degradation of system reliability. However, to our best knowledge, inference for partial shape constraints has not been recognized, although such interests can be plausible in practice.
@@
In this study, we propose the inferential tool to validate the partial shape constraints on functional regression coefficients in FoSR, along with two corresponding estimation approaches using kernel-smoothing and spline techniques. Suppose one is interested in the shape constraints on a sub-interval $\mathcal{I}=[a,b]\subset[0,1]$, where collected response trajectories span $[0,1]$. Under the existing tools for shape-constrained inference, one might consider using a part of functional observations restricted on $\mathcal{I}$ and apply the standard testing procedure. However, such an approach will cause a significant drop in sample sizes in the case of a bounded number of measurements on response trajectories, which is common in longitudinal studies, including our motivating example illustrated in Figure 1. The situation becomes much worse depending on the length of $\mathcal{I}$ or the sparsity of evaluation grids on it. Consequently, this naive approach would result in a severe deterioration in the power. To prevent this, we propose to borrow information from observations over the entire domain to perform the inference even for testing focused on specific subsets of the domain. It indeed prevents the significant drop in the power of the test while keeping the desirable level of type-I error, as demonstrated in the simulation studies of Section 3. The other contribution of our paper is in developing two partial shape-constrained regression estimators in parallel using the most widely used nonparametric smoothing techniques, kernel-smoothing and smoothing splines. The proposed unified testing tool applicable to both estimates would provide practical flexibility in a real application. We additionally derive asymptotic behaviors of the regression coefficient estimators with embedded partial shape constraints.
@@
The article is organized as follows. Section 2 presents the proposed method and results of the study. We introduce the partially shape-constrained FoSR models in Section 2.1. The estimation methods and theoretical findings are provided in Sections 2.2 and 2.3 that cover the kernel and spline estimators, respectively. In Section 2.4, we establish a unified inferential procedure for testing the partial shape constraints, encompassing kernel and spline approaches. Simulation results are reported in Section 3, and data applications are illustrated with two examples in Section 4. Technical details and proofs are provided in Supplementary Material.
2 Methodology and Theory
2.1 Testing partial shape constraints in FoSR models
Let $Y_{i}$ be the functional response coupled with a vector covariate $\mathbf{X}_{i}=(X_{i,1},\ldots,X_{i,p})^{\top}$ as
$Y_{i}(t)=\sum_{j=1}^{p}\beta_{j}(t)X_{i,j}+\varepsilon_{i}(t)\quad(t\in[0,1])$ (2.1)
@@ 2.1 Testing partial shape constraints in FoSR models
independently for $i=1,\ldots,n$, where $\boldsymbol{\beta}(t)=\big{(}\beta_{1}(t),\ldots,\beta_{p}(t)\big{)}^{\top}$ is a vector coefficient function and $\varepsilon_{i}(t)$ is a mean-zero stochastic process that models the regression error uncorrelated with $\mathbf{X}_{i}$, i.e., $\mathbb{E}(\varepsilon_{i}(t)|\mathbf{X}_{i})=0$. We assume that $X_{i,1},\ldots,X_{i,p}$ are linearly independent with positive probability, which allows the design matrix to include the intercept, e.g., $X_{i,1}=1$ for all $i=1,\ldots,n$, depending on the inferential interest. Despite the generality of the functional model (2.1), it is practically infeasible to observe the functional outcomes as the infinite-dimensional objects, and we assume that discrete evaluations $\mathbf{Y}_{i}=(Y_{i,1},\ldots,Y_{i,L_{i}})^{\top}$ with $Y_{i,\ell}=Y_{i}(T_{i,\ell})$ are only accessible for $\ell=1,\ldots,L_{i}$, where $T_{i,1},\ldots,T_{i,L_{i}}$ is a random sample of $T$ with a probability density function $\pi$ on $[0,1]$. Here, $L_{i}$ is an independent random integer that may or may not depend on the sample size $n$. More detailed conditions on $L_{i}$ required for kernel and regression spline methods are provided in the Supplementary Material S.1 and S.2. We denote the finite random sample as $\mathcal{X}_{n}=\{(\mathbf{Y}_{i},\mathbf{X}_{i}):i=1,\ldots,n\}$.
$H_{0,j}:\textrm{$\beta_{j}$ is $\mathcal{A}_{j}$-shaped on $\mathcal{I}_{j}$}$ (2.2)
@@
for $j\in J$. For example, if we hypothesize that $\beta_{j}$ is monotone increasing on $\mathcal{I}_{j}$, we read (2.2) as “$H_{0,j}:\textrm{$\beta_{j}$ is monotone increasing on $\mathcal{I}_{j}$}$.” Similarly, if one tests a partial convexity hypothesis, (2.2) can also be written as “$H_{0,j}:\textrm{$\beta_{j}$ is convex on $\mathcal{I}_{j}$}$.” In this paper, monotone and convex hypotheses will mainly be exemplified, yet one can extend our framework to general shape constraints, similarly as covered by many others in the literature [17, 18, 27, 28, 20, 25].
We then propose to reject the null hypothesis $H_{0}$ if the test statistic
is significantly large, where $\widehat{\boldsymbol{\beta}}(t)=\big{(}\hat{\beta}_{1}(t),\ldots,\hat{\beta}_{p}(t)\big{)}^{\top}$ and $\widetilde{\boldsymbol{\beta}}(t)=\big{(}\tilde{\beta}_{1}(t),\ldots,\tilde{\beta}_{p}(t)\big{)}^{\top}$ denote the estimates of coefficient function $\boldsymbol{\beta}(t)=\big{(}\beta_{1}(t),\ldots,\beta_{p}(t)\big{)}^{\top}$ under the null hypothesis $H_{0}$ and its general alternative $H_{1}$, respectively. While this $D_{n}$ is partly motivated by [22], a goodness-of-fit based test statistic for validating functional constraints on $\beta_{j}(t)$ via linear operator, similar types of the $L^{2}$-based test statistics were frequently employed in the literature for testing the nullity of functional difference for general scope [30, 36, 35]. Indeed, our proposed method corresponds to the global shape constraints if one sets $\mathcal{I}_{j}=[0,1]$. The following Sections 2.2 and 2.3 elaborate on how to obtain $\widehat{\boldsymbol{\beta}}(t)$ under kernel-smoothing and spline-smoothing approaches, respectively. Based on those estimators, we present a bootstrap test procedure in Section 2.4.
@@ 2.2 Estimation: Shape-constrained kernel-weighted least squares
We assume that the vector coefficient $\boldsymbol{\beta}(t)$ is as smooth as it allows local linear approximation $\boldsymbol{\beta}(u)\approx\boldsymbol{\beta}(t)+\dot{\boldsymbol{\beta}}(t)(u-t)$ for $u$ near $t$, where $\dot{\boldsymbol{\beta}}(t)=\big{(}\beta_{1}^{\prime}(t),\ldots,\beta_{p}^{\prime}(t)\big{)}^{\top}$ denotes the gradient of $\boldsymbol{\beta}(t)$. The local linear kernel estimator of $\mathbb{B}(t)=\big{[}\boldsymbol{\beta}(t),\dot{\boldsymbol{\beta}}(t)\big{]}\in\mathbb{R}^{p\times 2}$ is given by the unconstrained minimizer $\widetilde{\mathbb{B}}(t)$ of the following objective function.
$\begin{split}\mathcal{L}_{n}\big{(}\mathbb{B}(t)\big{)}&=\sum_{i=1}^{n}\frac{1}{L_{i}}\sum_{\ell=1}^{L_{i}}\Big{[}Y_{i,\ell}-\mathbf{X}_{i}^{\top}\mathbb{B}(t)\mathbf{Z}_{i,\ell}(t)\Big{]}^{2}K_{h}(T_{i,\ell}-t)\end{split}$ (2.4)

### v2

# Statistical inference for partially shape-constrained function-on-scalar linear regression models

@@ Abstract
Functional linear regression models are widely used to link functional/longitudinal outcomes with multiple scalar predictors, identifying time-varying covariate effects through regression coefficient functions. Beyond assessing statistical significance, characterizing the shapes of coefficient functions is crucial for drawing interpretable scientific conclusions. Existing studies on shape-constrained analysis primarily focus on global shapes, which require strict prior knowledge of functional relationships across the entire domain. This often leads to misspecified regression models due to a lack of prior information, making them impractical for real-world applications. To address this, a flexible framework is introduced to identify partial shapes in regression coefficient functions. The proposed partial shape-constrained analysis enables researchers to validate functional shapes within a targeted sub-domain, avoiding the misspecification of shape constraints outside the sub-domain of interest. The method also allows for testing different sub-domains for individual covariates and multiple partial shape constraints across composite sub-domains. Our framework supports both kernel- and spline-based estimation approaches, ensuring robust performance with flexibility in computational preference. Finite-sample experiments across various scenarios demonstrate that the proposed framework significantly outperforms the application of global shape constraints to partial domains in both estimation and inference procedures. The inferential tool particularly maintains the type I error rate at the nominal significance level and exhibits increasing power with larger sample sizes, confirming the consistency of the test procedure. The practicality of partial shape-constrained inference is demonstrated through two applications: a clinical trial on NeuroBloc for type A-resistant cervical dystonia and the National Institute of Mental Health Schizophrenia Study.
Keywords: Longitudinal/functional data, functional linear regression model, partial shape constraints, shape-constrained kernel least squares, shape-constrained regression spline
@@ 1 Introduction
As function-valued data acquisition becomes increasingly common, there has been a significant amount of work devoted to functional data analysis (FDA). The high- or infinite-dimensional structure of functional data offers many analytical opportunities across various fields, from comprehensive data exploration to modeling with its rich source of information; see, for example, [31], [11], [16], and [19]. Among advances in FDA, functional linear regression has received attention to validate and identify relationships between functional responses and scalar or functional predictors [29, 40, 14]. Specifically, statistical analysis with function-on-scalar regression (FoSR) models has been making implications in various applications [33] by capturing the dynamic association between functional responses and scalar covariates, varying over the domain, through coefficient function estimation and inferential procedures.
Let $Y_{i}$ be the functional response associated with multiple covariates $X_{i,1},\ldots,X_{i,p}$, then FoSR model is written as
$Y_{i}(t)=\sum_{j=1}^{p}\beta_{j}(t)X_{i,j}+\varepsilon_{i}(t)\quad(t\in[0,1]),$ (1.1)
@@
for each $i=1,\ldots,n$ independently, where $\beta_{1}(t),\ldots,\beta_{p}(t)$ are coefficient functions and $\varepsilon_{i}(t)$ is a mean-zero stochastic process such that $\mathrm{E}(\varepsilon_{i}(t)|\mathbf{X}_{i})=0$ with $\mathbf{X}_{i}=(X_{i,1},\ldots,X_{i,p})^{\top}$. For technical convenience, we assume that the $n\times p$ design matrix, consisting of row vectors $\mathbf{X}_{1}^{\top},\ldots,\mathbf{X}_{n}^{\top}$, has full column rank, and we allow the design matrix to include the intercept, e.g., $X_{i,1}=1$ for all $i=1,\ldots,n$, depending on the inferential interest. Then, existing studies on inferential problems in FoSR generally focus on developing inferential tools to test the null hypothesis $H_{0}:\beta_{j}=0$, for $j=1,\ldots,p,$ assessing the significance of the coefficients [38, 44, 43] However, besides finding the statistical evidence of a non-null covariate effect on response trajectories over the domain, validating the shape of functional regression coefficients, such as monotonicity, convexity, or concavity, over a specific subset of the domain can be crucial for domain experts to allow tangible interpretation in practice. Once the functional shape is validated, one should incorporate the corresponding conditions to the model estimates to avoid potential biases. Or, in practical estimation problems, shape restrictions on functional regression coefficients over specific sub-intervals of the domain can be known as prior knowledge.
@@
Figure 1: The left panel illustrates the individual trajectories of disease severity for a subset of patients in the treatment and placebo groups. The average disease severity for each group (solid line) shows an overall decrease, with the treatment group experiencing a rapid drop in the first few weeks, followed by a continual reduction at a seemingly similar rate for both groups. As shown in the right panel, Ghosal et al. (2023) previously validated this observation using globally monotone shape-constrained estimates ( black dotted line), which provide more explicit and decision-making interpretations than unconstrained estimates ( black dashed line). However, the global shape-constrained inference may be misleading by ignoring the rebound after Week 5, as indicated by the unconstrained estimates. Our proposed method (blue solid line) with partial shape constraints refines the conclusion, suggesting that the duration of drug action may effectively extend up to Week 3, with sustained efficacy thereafter.
Our motivating example is the functional data analysis approach to demonstrating the treatment efficacy of drug use during the period of clinical trials in the National Institute of Mental Health Schizophrenia Collaborative Study. The previous studies, including [1] and [13], showed a significant continual drop in disease severity for the treatment group, i.e., improving drug efficacy over weeks through the shape-constraint test on the drug effect coefficient function. While the study design and models are provided in Section 4.2, the left panel of Figure 1 displays the collected longitudinal disease severity measures surveyed from the initial treatment (week 0) to week 6 from randomly selected 30 patients of placebo and treatment cohorts. However, refined conclusions over specific periods can be of practical interest to practitioners, such as when the maximum drug efficacy is achieved or whether we can conclude significant improvement in drug effectiveness even during the later weeks of trials. As we shall see in Section 4.2, our proposed partial inference method concludes the significant monotone decrease in disease severity owing to the drug treatment over weeks 0–3, followed by the consistent duration of such efficacy for the remainder weeks, illustrated with the blue solid line in the right panel of Figure 1. This is somewhat distinct from constrained estimates under the condition of monotone decrease over the entire domain, the conclusion of [13], and unconstrained smooth estimates, which are displayed with the black dotted and dashed lines, respectively.
@@
The recent literature on functional data analysis and its applications also pays attention to shape-restricted estimation and inference in functional regression models. For the inference on shape-constrained regression coefficients, [6] proposed the goodness-of-fit test based on empirical processes projected to the space with shape constraints, while [30] addresses a similar problem under the null hypothesis with linear operator constraints. For estimation, [4] extended the shape-constrained kernel smoothing to the functional and longitudinal data framework, and [13] employed the Bernstein polynomials for shape-constrained estimation in functional regression with one of the data application results from [13] being illustrated in Figure 1. To name of few other applications, in economics, motivated by the pioneering work from Slutsky [39] recognizing the necessity of shape-restricted estimation and inference, [5] elaborated theoretical and application works in econometric research. In public health research, the analysis of growth charts under the monotone increasing shape restrictions has provided a crucial clinical tool for growth screening during infancy, childhood, and adolescence [20] or such growth chart helps health providers assess monotone increasing growth patterns against age-specific percentile curves [9]. The reliability engineering community also paid attention to this topic for bathtub-shaped function estimation in assessing the degradation of system reliability. However, to our best knowledge, inference for partial shape constraints has not been recognized, although such interests can be plausible in practice.
Figure 2: An illustration of partial shape-constrained analysis: When focusing on a sub-domain of inferential interest, existing tools restrict functional observations to that domain, introducing significant bias from boundary issues in function estimation and substantially reducing the power of the analysis due to the smaller sample size. We propose a framework that leverages all available observations while applying partial shape constraints, preserving the same statistical power as a complete data analysis.
@@
In this study, we propose the inferential tool to validate the partial shape constraints on functional regression coefficients in FoSR, along with two corresponding estimation approaches using kernel smoothing and spline techniques, respectively. Suppose one is interested in the shape constraints on a sub-interval $\mathcal{I}=[a,b]\subset[0,1]$, where collected response trajectories span $[0,1]$. Under the existing tools for shape-constrained inference, one might consider using a part of functional observations restricted on $\mathcal{I}$ and apply the standard testing procedure. However, such an approach will cause a significant drop in sample sizes in the case of a bounded number of measurements on response trajectories, which is common in longitudinal studies, as illustrated by the toy example in Figure 2. The situation becomes much worse depending on the length of $\mathcal{I}$ or the sparsity of evaluation grids on it. Consequently, this naive approach would result in a severe deterioration in the power. We thus propose to borrow information from observations over the entire domain to perform the inference even for testing focused on specific subsets of the domain.
Our article has the following major contributions. The proposed method mitigates a significant drop in the power of the test. Even when applied to a partial domain, it maintains the desirable level of type-I error, as demonstrated in the simulation studies of Section 3. Another key contribution is the development of two partial shape-constrained regression estimators in parallel, using the two most widely used nonparametric smoothing techniques: kernel- and spline-based smoothing. The proposed unified testing tool applicable to both estimates would provide practical flexibility in a real application. Additionally, we derive asymptotic behaviors of the regression coefficient estimators with embedded partial shape constraints to ensure consistent behaviors under large samples. In this paper, monotone and convex hypotheses will mainly be exemplified. Yet, one can extend our framework to general shape constraints, including partial positivity, partial boundedness, and their combinations, similarly as covered by many others in the literature [23, 24, 36, 27, 34].
@@
The article is organized as follows. Section 2 presents the proposed method and results of the study. The estimation methods and theoretical findings are provided in Sections 2.2 and 2.3 that cover the kernel and spline estimators, respectively. In Section 2.4, we establish a unified inferential procedure for testing the partial shape constraints, encompassing kernel and spline approaches. Simulation results are reported in Section 3, and data applications are illustrated with two examples in Section 4. Technical details and proofs are provided in the online Supplementary Material.
2 Methodology
2.1 Problem setup
@@
We begin by specifying assumptions on discretized evaluations of functional responses and corresponding notations used, followed by elaborated null hypotheses considered in our study. While the FoSR model, presented in (1.1), is generally formulated over a continuous domain, it is practically infeasible to observe the functional outcomes as infinite-dimensional objects. We thus assume that discrete evaluations $\mathbf{Y}_{i}=(Y_{i,1},\ldots,Y_{i,L_{i}})^{\top}$ with $Y_{i,\ell}=Y_{i}(T_{i,\ell})$ are only accessible for $\ell=1,\ldots,L_{i}$, where $T_{i,1},\ldots,T_{i,L_{i}}$ is a random sample of $T$ with a probability density function $\pi$ on $[0,1]$. Here, $L_{i}$ is an independent random integer that may or may not depend on the sample size $n$. More detailed conditions on $L_{i}$ required for kernel and regression spline methods are provided in Sections A and B of the online Supplementary Material. We denote the finite random sample as $\mathcal{X}_{n}=\{(\mathbf{Y}_{i},\mathbf{X}_{i}):i=1,\ldots,n\}$.
@@
Then we want to estimate regression coefficient functions subject to shape constraints restricted on pre-specified sub-intervals, denoted by $\mathcal{I}_{j}\subset[0,1]$ for $j\in J$ with $J\subset\{1,\ldots,p\}$. For example, if we hypothesize $H_{0,j}$: $\beta_{j}$ is partially monotone increasing on $\mathcal{I}_{j}$ or $H_{0,k}$: $\beta_{k}$ is partially convex on $\mathcal{I}_{k}$ for some $j,k\in J$, where $j$ and $k$ may or may not be identical, we want to estimate the model (1.1) under the null hypothesis so that we can identify the most likely partial shapes. In this paper, monotone and convex hypotheses will mainly be exemplified. Yet, one can extend our framework to general shape constraints, including partial positivity, partial boundedness, and their combinations, similarly as covered by many others in the literature [23, 24, 36, 27, 34].
@@ 2.2 Estimation: Shape-constrained kernel-weighted least squares
We assume that the vector coefficient $\boldsymbol{\beta}(t)=\big{(}\beta_{1}(t),\ldots,\beta_{p}(t)\big{)}^{\top}$ is as smooth as it allows local linear approximation $\boldsymbol{\beta}(u)\approx\boldsymbol{\beta}(t)+\dot{\boldsymbol{\beta}}(t)(u-t)$ for $u$ near $t$, where $\dot{\boldsymbol{\beta}}(t)=\big{(}\beta_{1}^{\prime}(t),\ldots,\beta_{p}^{\prime}(t)\big{)}^{\top}$ denotes the gradient of $\boldsymbol{\beta}(t)$. The local linear kernel estimator of $\mathbb{B}(t)=\big{[}\boldsymbol{\beta}(t),\dot{\boldsymbol{\beta}}(t)\big{]}\in\mathbb{R}^{p\times 2}$ is given by the unconstrained minimizer $\widetilde{\mathbb{B}}(t)$ of the following objective function.
$\begin{split}\mathcal{L}_{n}\big{(}\mathbb{B}(t)\big{)}&=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{L_{i}}\sum_{\ell=1}^{L_{i}}\Big{[}Y_{i,\ell}-\mathbf{X}_{i}^{\top}\mathbb{B}(t)\mathbf{Z}_{i,\ell}(t)\Big{]}^{2}K_{h}(T_{i,\ell}-t),\end{split}$ (2.1)
@@
We list the technical conditions for theoretical results. For the minimum and maximum measurement frequencies $\lambda_{n}=\min_{1\leq i\leq n}L_{i}$ and $\Lambda_{n}=\max_{1\leq i\leq n}L_{i}$, let $N=n\lambda_{n}$.
  1. $(K0)$
    The probability density function $\pi$ of $T$ is continuously differentiable and strictly positive.
  2. $(K1)$
  3. $(K2)$
    $E\|\varepsilon\|_{\infty}^{k}<\infty$ for some $k>2$.
  4. $(K3)$
    $h\asymp N^{-a}$ for some $0<a<1-2/k$.
  5. $(K4)$
    $\lambda_{n}=O(1)$ and $\Lambda_{n}=O(N^{b}/\log N)^{k/2}$ for some $0<b<1-(2/k)-a$.